If a simple pendulum has significant amplitude (up to a factor of 1/e of original) only in the period between t = Os to t = ts, then t may be called the average life of the pendulum. When the spherical bob of the pendulum suffers a retardation (due to viscous drag) proportional to its velocity, with 'b' as the constant of proportionality, the average life time of the pendulum is (assuming damping is small) in seconds: (1) 0.693 b (2) b (3) 1 b (4) 2 b

Grade:11

1 Answers

shruti
10 years ago
equation of motion is: dv/dt = -kx/m -bv amplitude is A = Ao exp(-bt/2) t= ts A = Ao/e so, t = 2/b